# MemBuilder: Reinforcing LLMs for Long-Term Memory Construction via Attributed Dense Rewards

### v2

@@ Abstract
Zhiyu Shen1thanks: Equal contribution.thanks: Work done during internship at Tencent Inc., Ziming Wu2∗thanks: Corresponding author., Fuming Lai2, Shaobing Lian2, Yanghui Rao1 1School of Computer Science and Engineering, Sun Yat-sen University. 2Tencent Inc. shenzhy23@mail2.sysu.edu.cn {jimmyzmwu, fuminglai, lokilian}@tencent.com raoyangh@mail.sysu.edu.cn
@@ 1 Introduction
Memory-augmented frameworks have emerged as a promising approach for maintaining consistency in long-term dialogues, which must track evolving contexts and historical states over extended timelines. While Retrieval-Augmented Generation (RAG) facilitates access to external knowledge, it treats retrieval units as independent, static chunks—failing to capture how information evolves or which historical facts have been superseded (Liu et al., 2024; Gao et al., 2024). Memory-augmented frameworks address this by decomposing information prior to storage: events receive independent timestamps and semantic concepts are structured into discrete units. This shifts the computational burden from processing entangled tokens during inference to retrieving precise, “pre-digested" fragments. Recent implementations like Mem0 Chhikara et al. (2025), MIRIX Wang and Chen (2025), and MemGPT Packer et al. (2023) exemplify this approach, constructing external memory that evolves with each interaction. However, these systems rely largely on fixed prompting templates and expensive closed-source models, operating in an “open loop" without feedback on whether the constructed memories actually benefit downstream tasks. This raises a critical question: Can we instead train a model to perform memory construction through direct supervision?
@@
We present MemBuilder, a framework for training models to construct long-term memory with attributed dense rewards. Our architecture utilizes a multi-dimensional memory design, comprising Core, Episodic, Semantic and Procedural components, and trains a single, lightweight 4B model to manage all four components. Our approach introduces two key technical contributions to address the limitations of existing RL-based memory construction methods (Figure 1). First, we employ dense session-level rewards. Unlike traditional methods that assign a single reward after processing all sessions, we leverage synthetic session-level question generation to provide immediate feedback after each session’s memory operations. Second, we introduce contribution-aware gradient weighting to resolve multi-dimensional memory attribution. Since all four memory components share a common reward, their individual contributions remain ambiguous. Our mechanism addresses this by scaling gradient updates based on the downstream utility of the constructed memories; specifically, the gradient impact of a memory operation is proportional to the usage of its corresponding memory component during retrieval. We integrate these techniques into Attributed Dense Rewards Policy Optimization (ADRPO). Extensive experiments conducted on different benchmarks demonstrate the effectiveness of our proposed framework. Specifically, MemBuilder achieves 84.23% on LoCoMo, surpassing the baselines including Claude 4.5 Sonnet under the same setting.
@@ 2.2 Prompting-Based Memory Frameworks
2.3 Learning-Based Approaches for Memory Construction
@@ 3.2 Multi-Dimensional Memory Architecture
  • Core Memory $\mathcal{M}^{\text{core}}$: A fixed-size block storing persistent user profile information including identity, preferences, and key relationships. This memory is always included in the context during question answering. Detailed prompt templates are provided in Appendix H.
  • Episodic Memory $\mathcal{M}^{\text{epi}}$: Time-stamped event records capturing what happened and when. Each entry follows the format “YYYY-MM-DD: Event summary | Details”, enabling temporal reasoning.
  • Semantic Memory $\mathcal{M}^{\text{sem}}$: Factual knowledge about entities in the user’s life, such as people, places, and user-specific concepts. Common knowledge is explicitly excluded to avoid redundancy.
  • Procedural Memory $\mathcal{M}^{\text{proc}}$: Step-by-step processes, routines, and workflows mentioned in conversations, such as the user’s morning routine or problem-solving approach.
Given a new conversation session, "all four memory types are processed simultaneously, each extracting memories according to its specialized perspective. Core Memory is maintained as a fixed block with automatic compression when capacity is exceeded. The other three memory types are stored in a vector database and retrieved via semantic similarity during question answering.
@@
$\mathcal{A}^{\text{core}}$ $=\{\textsc{Append},\textsc{Replace},\textsc{Rewrite}\}$ (1)
$\mathcal{A}^{\text{epi}}$ $=\{\textsc{Add},\textsc{Update},\textsc{Merge}\}$ (2)
$\mathcal{A}^{\text{sem}}$ $=\{\textsc{Add},\textsc{Update},\textsc{Skip}\}$ (3)
$\mathcal{A}^{\text{proc}}$ $=\{\textsc{Add},\textsc{Update}\}$ (4)
@@ 3.3 Supervised Fine-Tuning
While multi-dimensional memory architectures can offer finer-grained control (Wang and Chen, 2025), lightweight models, such as Qwen3-4B employed in our framework, often struggle with direct multi-dimensional memory construction, frequently producing invalid actions that impede effective RL exploration. To address this cold-start problem, we collect expert trajectories $\{(\mathit{State}_{t},a_{t})\}_{t=1}^{n}$ using Claude 4.5 Sonnet. This phase stabilizes the model’s output format, providing a viable baseline for subsequent training. However, SFT primarily focuses on behavioral cloning with limited exploration capability; we further employ RL to maximize the model’s ultimate utility in QA contexts.
@@ 3.4.1 Dense Session-Level Rewards via Synthetic Session-level QA
We address this through synthetic session-level QA that evaluates memory quality at each step. Before the RL training, for each session $\tau$, we retrieve the top-$k$ memories from $\mathcal{M}_{\tau-1}$ most similar to the session $s_{\tau}$, and let an expert model with $(s_{\tau},\mathcal{M}_{\text{retrieved}})$ generate $J$ question-answer pairs $\{(q_{j},\mathit{ans}_{j})\}_{j=1}^{J}$ targeting information in $s_{\tau}$ or its connections to retrieved memoey. Questions span three types: single-session (testing current session retention), multi-session (requiring cross-session aggregation), and temporal-reasoning (involving time-based inference).
@@
$\ell^{(\text{core})}=\begin{cases}0&\text{if }\Delta_{\text{core}}\leq\theta_{\min}\\[3.0pt] \frac{\Delta_{\text{core}}-\theta_{\min}}{\theta_{\max}-\theta_{\min}}&\text{if }\Delta_{\text{core}}\in(\theta_{\min},\theta_{\max})\\[3.0pt] 1&\text{if }\Delta_{\text{core}}\geq\theta_{\max}\end{cases}$ (7)
@@
$\ell^{(m)}=\begin{cases}0&\text{if }\Delta<\delta\text{ or }\rho\in[\gamma_{l},\gamma_{u}]\\[3.0pt] \frac{\rho-\gamma_{u}}{\gamma_{\max}-\gamma_{u}}&\text{if }\rho\in(\gamma_{u},\gamma_{\max}]\\[3.0pt] \frac{\gamma_{l}-\rho}{\gamma_{l}-\gamma_{\min}}&\text{if }\rho\in[\gamma_{\min},\gamma_{l})\\[3.0pt] 1&\text{otherwise}\end{cases}$ (8)
@@ 3.4.2 Contribution-Aware Gradient Weighting
Within each rollout, actions operating on the four memory dimensions contribute to a shared memory bank and receive a global reward. However, the functional impact on downstream QA performance varies significantly across memory types; for instance, Episodic Memories may be frequently retrieved while Procedural Memories remain unused. To account for these discrepancies, we dynamically amplify gradient updates based on each component’s downstream utility.
During QA evaluation, we record retrieval counts $h^{(m)}$ for each memory type $m\in\{\mathrm{epi},\mathrm{sem},\mathrm{proc}\}$ across all questions. The dominant contributing type is:
$d=\arg\max_{m\in\{\mathrm{epi},\mathrm{sem},\mathrm{proc}\}}h^{(m)}$ (9)
@@
$w^{(m)}=\begin{cases}\alpha&\text{if }m=d\\ 1&\text{otherwise}\end{cases}$ (10)
where $\alpha>1$ amplifies updates for the dominant contributor. Core Memory, which is always included in the context rather than retrieved, receives a fixed weight $w^{(\mathrm{core})}=1$. This mechanism ensures that memory types whose entries directly contributed to successful QA receive proportionally stronger reinforcement.
@@ 3.4.3 Training Objective
$\mathcal{J}(\theta)=\mathbb{E}\Bigg[\sum_{m}\frac{1}{|a_{i}^{(m)}|}\sum_{k=1}^{|a_{i}^{(m)}|}\min\Big(w^{(m)}\rho_{i,k}^{(m)}A_{i},\\ \text{clip}(\rho_{i,k}^{(m)},1\!-\!\epsilon,1\!+\!\epsilon)A_{i}\Big)\Bigg]-\beta\cdot D_{\mathrm{KL}}(\pi_{\theta}\|\pi_{\mathrm{ref}})$ (12)
where $\rho_{i,k}^{(m)}=\pi_{\theta}/\pi_{\mathrm{ref}}$ is the importance ratio for the $k$-th token of memory type $m$’s output. The contribution-aware weights $w^{(m)}$ scale the unclipped term, encouraging larger updates for high-impact components while preserving the clipping mechanism for stability.
@@ 4.1 Experimental Setup
We use Qwen3-4B-Instruct-2507 as our base model, with SFT trajectories collected using Claude 4.5 Sonnet. GPT-4.1 serves as the LLM judge for evaluation. All retrieval uses text-embedding-3-small. To isolate memory construction quality, we fix the answer model to Claude 4.5 Sonnet unless otherwise specified. Full details are in Appendix B and D. Detailed configuration including embedding settings and action formats is in Appendix C. A detailed cost breakdown is provided in Appendix F.
@@ 4.3.1 Effect of Gradient Weighting
To investigate the effect of contribution-aware gradient weighting (Section 3.4.2), we vary the weighting coefficient $\alpha$ that amplifies updates for the dominant contributing memory type. We conduct this ablation on a reduced training set for efficiency. As shown in Figure 4, performance improves as $\alpha$ increases from 1 (no weighting, 82.82%) to 4 (84.04%), confirming that attributing credit to high-contribution memory types enhances the final model performance. However, excessively large $\alpha$ values degrade performance due to gradient imbalance among memory types, with the optimal value at $\alpha=4$.
@@ 4.3.3 Answer Model Generalization
To evaluate whether the constructed memory generalizes across different answer models, we fix the memory construction model to our trained Qwen3-4B and vary the answer model. Table 2 shows that our memory maintains high quality across answer models of varying capabilities.
@@ A.1 Benchmark Datasets
A.2 Training Data Construction
We use LongMemEval as our sole training source. From the 500 available dialogues, we sample 50 dialogues for SFT trajectory collection and a separate 50 dialogues for RL dataset construction. All other benchmarks (LoCoMo and PerLTQA) serve as out-of-distribution test sets to evaluate generalization.
SFT Dataset.
The 50 SFT dialogues comprise approximately 2,400 sessions. Since each session requires memory operations from all four memory types (Core, Episodic, Semantic, Procedural) and we train each memory type’s output as a separate example, this yields 9,600 training samples. Each sample pairs an input (retrieved memories + current session) with a single type of memory operations in JSON format. Based on logged data, the average input length is approximately 6,000 tokens and the average output length is approximately 780 tokens, resulting in a total of approximately 65M tokens for SFT training.
RL Dataset.
The 50 RL dialogues comprise approximately 2,400 sessions. For each session, we generate 5 synthetic QA pairs for dense reward computation, resulting in 12,000 QA pairs in total. During GRPO training with 8 rollouts per session and 5 epochs, this produces 96,000 session-rollouts for policy optimization.
Appendix B Training Pipeline Details
B.1 Expert Trajectory Collection
We collect expert memory construction trajectories using Claude 4.5 Sonnet. For each dialogue, we process sessions sequentially: at step $k$, the model receives the new session $s_{k}$ along with relevant memories retrieved from the step k1k-1 memory bank, then generates memory operations for all four memory types. After each step, we compute embeddings for newly created memories and update the vector database. This produces complete trajectories of memory states $\{\mathcal{M}_{0},\mathcal{M}_{1},\ldots,\mathcal{M}_{n}\}$ for each dialogue.
B.2 RL Dataset Construction
The RL dataset construction involves three components:
Input Preparation.
For each session $k$, we construct the policy model’s input by retrieving the top-20 most relevant memories from the step k1k-1 vector database, concatenated with the new session $s_{k}$. This mirrors the inference-time setup where the model must decide what to extract given limited context about historical memories.
Synthetic Question Generation.
To enable dense session-level rewards, we generate 5 question-answer pairs for each session using Claude 4.5 Opus. The generation model receives the new session $s_{k}$ along with the top-20 most similar memories retrieved from the step k1k-1 memory bank. This retrieval provides context about historical information, allowing the model to generate questions that test not only the understanding of the new session content but also its connections to prior history when relevant relationships exist. The model is instructed to create questions spanning factual recall, temporal reasoning, and inference tasks.
Reward Computation.
During the RL training, each session undergoes 8 rollouts, producing 8 candidate memory banks $\{\mathcal{M}_{k}^{(1)},\ldots,\mathcal{M}_{k}^{(8)}\}$. For each candidate, GPT-4.1-mini answers the pre-generated questions by retrieving from the candidate memory bank, and following prior work (Chhikara et al., 2025; Wang and Chen, 2025), and an LLM judge (GPT-4.1-mini) evaluates answer correctness. The average accuracy across 5 questions yields the task reward for each rollout. Note that within each rollout, all four memory types share the same task reward, which is then differentiated through contribution-aware gradient weighting (Section 3.4.2).
B.3 SFT Training

### v3

@@ Abstract
Zhiyu Shen1thanks: Equal contribution.thanks: Work done during internship at Tencent Inc., Ziming Wu2∗thanks: Corresponding author., Fuming Lai2, Shaobing Lian2, Yanghui Rao1 1School of Computer Science and Engineering, Sun Yat-sen University, Guangzhou, China 2Tencent Inc. shenzhy23@mail2.sysu.edu.cn, raoyangh@mail.sysu.edu.cn {jimmyzmwu, fuminglai, lokilian}@tencent.com
@@ 1 Introduction
Memory-augmented frameworks have emerged as a promising approach for maintaining consistency in long-term dialogues, which must track evolving contexts and historical states over extended timelines. While Retrieval-Augmented Generation (RAG) facilitates access to external knowledge, it treats retrieval units as independent, static chunks—failing to capture how information evolves or which historical facts have been superseded (Liu et al., 2024; Gao et al., 2024). Memory-augmented frameworks address this by decomposing information prior to storage: events receive independent timestamps and semantic concepts are structured into discrete units. This shifts the computational burden from processing entangled tokens during inference to retrieving precise, “pre-digested" fragments. Recent implementations like Mem0 Chhikara et al. (2025), MIRIX Wang and Chen (2025), and MemGPT Packer et al. (2023) exemplify this approach, constructing external memory that evolves with each interaction. However, these prompting-based systems face three key limitations: they operate in an “open loop” without feedback on whether the constructed memories benefit downstream tasks; their performance degrades sharply when substituting less capable models; and their reliance on expensive closed-source models for the token-intensive memory construction phase makes deployment cost prohibitive at scale. This raises a critical question: Can we instead train a model to perform memory construction through direct supervision?
@@
We present MemBuilder, a framework for training models to construct long-term memory with attributed dense rewards. Our architecture utilizes a multi-dimensional memory design, comprising Core, Episodic, Semantic and Procedural components, and trains a single, lightweight 4B model to manage all four components. Our approach introduces two key technical contributions to address the limitations of existing RL-based memory construction methods (Figure 1). First, we employ dense session-level rewards. Unlike traditional methods that assign a single reward after processing all sessions, we leverage synthetic session-level question generation to provide immediate feedback after each session’s memory operations. Second, we introduce contribution-aware gradient weighting to resolve multi-dimensional memory attribution. Since all four memory components share a common reward, their individual contributions remain ambiguous. Our mechanism addresses this by scaling gradient updates based on the downstream utility of the constructed memories; specifically, the gradient impact of a memory operation is proportional to the usage of its corresponding memory component during retrieval. We integrate these techniques into Attributed Dense Rewards Policy Optimization (ADRPO). Extensive experiments conducted on different benchmarks demonstrate the effectiveness of our proposed framework. Specifically, MemBuilder achieves 84.23% on LoCoMo, surpassing the baselines including Claude 4.5 Sonnet under the same setting. Notably, our model is trained exclusively on LongMemEval yet generalizes effectively to out-of-distribution benchmarks with different dialogue structures and question types.
@@ 2.2 Prompting-Based Memory Frameworks
2.3 Training-Based Approaches for Memory Construction
@@ 3.2 Multi-Dimensional Memory Architecture
  • Core Memory $\mathcal{M}^{\textrm{core}}$: A fixed-size block storing persistent user profile information including identity, preferences, and key relationships. This memory is always included in the context during question answering. Detailed prompt templates are provided in Appendix I.
  • Episodic Memory $\mathcal{M}^{\textrm{epi}}$: Time-stamped event records capturing what happened and when. Each entry follows the format “YYYY-MM-DD: Event summary | Details”, enabling temporal reasoning.
  • Semantic Memory $\mathcal{M}^{\textrm{sem}}$: Factual knowledge about entities in the user’s life, such as people, places, and user-specific concepts. Common knowledge is explicitly excluded to avoid redundancy.
  • Procedural Memory $\mathcal{M}^{\textrm{proc}}$: Step-by-step processes, routines, and workflows mentioned in conversations, such as the user’s morning routine or problem-solving approach.
Given a new conversation session, all four memory types are processed simultaneously, each extracting memories according to its specialized perspective. Core Memory is maintained as a fixed block with automatic compression when capacity is exceeded. The other three memory types are stored in a vector database and retrieved via semantic similarity during question answering.
@@
$\mathcal{A}^{\textrm{core}}$ $=\{\textsc{Append},\textsc{Replace},\textsc{Rewrite}\}$ (1)
$\mathcal{A}^{\textrm{epi}}$ $=\{\textsc{Add},\textsc{Update},\textsc{Merge}\}$ (2)
$\mathcal{A}^{\textrm{sem}}$ $=\{\textsc{Add},\textsc{Update},\textsc{Skip}\}$ (3)
$\mathcal{A}^{\textrm{proc}}$ $=\{\textsc{Add},\textsc{Update}\}$ (4)
@@ 3.3 Supervised Fine-Tuning
While multi-dimensional memory architectures can offer finer-grained control (Wang and Chen, 2025), lightweight models, such as Qwen3-4B employed in our framework, often struggle with direct multi-dimensional memory construction, frequently producing invalid actions—such as malformed JSON structures, missing required fields (e.g., episodic entries without timestamps), or misuse of undefined operation types—that impede effective RL exploration. To address this cold-start problem, we collect expert trajectories $\{(\mathit{State}_{t},a_{t})\}_{t=1}^{n}$ using Claude 4.5 Sonnet. This phase stabilizes the model’s output format, providing a viable baseline for subsequent training. However, SFT, as a form of imitation learning, trains the model to replicate expert memory operations but does not directly optimize for downstream QA accuracy; we further employ RL to maximize the model’s ultimate utility in QA contexts.
@@ 3.4.1 Dense Session-Level Rewards via Synthetic Session-level QA
We address this through synthetic session-level QA that evaluates memory quality at each step. Before the RL training, for each session $\tau$, we retrieve the top-$k$ memories from $\mathcal{M}_{\tau-1}$ most similar to the session $s_{\tau}$, and let an expert model with $(s_{\tau},\mathcal{M}_{\textrm{retrieved}})$ generate $J$ question-answer pairs $\{(q_{j},\mathit{ans}_{j})\}_{j=1}^{J}$ targeting information in $s_{\tau}$ or its connections to retrieved memory.
@@
$\ell^{(\text{core})}=\begin{cases}0&\textrm{if }\Delta_{\text{core}}\leq\theta_{\min}\\[3.0pt] \frac{\Delta_{\text{core}}-\theta_{\min}}{\theta_{\max}-\theta_{\min}}&\textrm{if }\Delta_{\text{core}}\in(\theta_{\min},\theta_{\max})\\[3.0pt] 1&\textrm{if }\Delta_{\text{core}}\geq\theta_{\max}\end{cases}$ (7)
@@
$\ell^{(m)}=\begin{cases}0&\textrm{if }\Delta<\delta\textrm{ or }\rho\in[\gamma_{l},\gamma_{u}]\\[3.0pt] \frac{\rho-\gamma_{u}}{\gamma_{\max}-\gamma_{u}}&\textrm{if }\rho\in(\gamma_{u},\gamma_{\max}]\\[3.0pt] \frac{\gamma_{l}-\rho}{\gamma_{l}-\gamma_{\min}}&\textrm{if }\rho\in[\gamma_{\min},\gamma_{l})\\[3.0pt] 1&\textrm{otherwise}\end{cases}$ (8)
@@ 3.4.2 Contribution-Aware Gradient Weighting
Within each rollout, actions operating on the four memory dimensions contribute to a shared memory bank and receive a global reward. However, the functional impact on downstream QA performance varies significantly across memory types; for instance, Episodic Memories may be frequently retrieved while Procedural Memories remain unused. To account for these discrepancies, we dynamically amplify gradient updates based on each component’s downstream utility. Since vanilla GRPO computes a single advantage per rollout and applies it uniformly to all components, it cannot distinguish their individual contributions—a challenge analogous to credit assignment in cooperative multi-agent settings (Foerster et al., 2018; Li et al., 2022). Drawing on prior work that scales gradient updates by downstream utility to resolve such attribution ambiguity (Li et al., 2022; Ren et al., 2018), we dynamically amplify gradient updates based on each component’s retrieval-based contribution.
During QA evaluation, we record retrieval counts $h^{(m)}$ for each memory type $m\in\{\textrm{epi},\textrm{sem},\textrm{proc}\}$ across all questions. The dominant contributing type is:
$d=\arg\max_{m\in\{\textrm{epi},\textrm{sem},\textrm{proc}\}}h^{(m)}$ (9)
@@
$w^{(m)}=\begin{cases}\alpha&\textrm{if }m=d\\ 1&\textrm{otherwise}\end{cases}$ (10)
where $\alpha>1$ amplifies updates for the dominant contributor. Core Memory, which is always included in the context rather than retrieved, receives a fixed weight $w^{(\textrm{core})}=1$. This mechanism ensures that memory types whose entries directly contributed to successful QA receive proportionally stronger reinforcement.
@@ 3.4.3 Training Objective
$\mathcal{J}(\theta)=\mathbb{E}\Bigg[\sum_{m}\frac{1}{|a_{i}^{(m)}|}\sum_{k=1}^{|a_{i}^{(m)}|}\min\Big(w^{(m)}\rho_{i,k}^{(m)}A_{i},\\ w^{(m)}\cdot\text{clip}(\rho_{i,k}^{(m)},1\!-\!\epsilon,1\!+\!\epsilon)A_{i}\Big)\Bigg]-\beta\cdot D_{\textrm{KL}}(\pi_{\theta}\|\pi_{\textrm{ref}})$ (12)
where $\rho_{i,k}^{(m)}=\pi_{\theta}/\pi_{\textrm{ref}}$ is the importance ratio for the $k$-th token of memory type $m$’s output. The contribution-aware weights $w^{(m)}$ scale both terms within the $\min$ operator, uniformly amplifying the objective for high-impact components while preserving the clipping mechanism for stability.
@@ 4.1 Experimental Setup
We use Qwen3-4B-Instruct-2507 as our base model. GPT-4.1 serves as the LLM judge for evaluation. All retrieval uses text-embedding-3-small. To isolate memory construction quality, we fix the answer model to Claude 4.5 Sonnet unless otherwise specified. Training hyperparameters are provided in Appendix B and baseline details in Appendix D. Detailed configuration including embedding settings and action formats is in Appendix C. A detailed cost breakdown is provided in Appendix F.
4.2 Training Data Construction
We use LongMemEval as our sole training source, sampling 50 dialogues for SFT and a separate 50 for RL. All other benchmarks serve as OOD test sets.
SFT Dataset.
The 50 SFT dialogues comprise approximately 2,400 sessions. We collect expert trajectories using Claude 4.5 Sonnet by processing sessions sequentially: at step $k$, the model receives the new session $s_{k}$ along with the top-20 memories retrieved from the step k1k{-}1 memory bank, then generates memory operations for all four types. After each step, newly created memories are embedded and added to the vector database. Since each session produces four memory type outputs trained as separate examples, this yields 9,600 training samples with an average input length of 6,000 tokens and output length of 780 tokens.
RL Dataset.
The 50 RL dialogues similarly comprise approximately 2,400 sessions. For each session $\tau$, we construct the policy model’s input by retrieving the top-20 most relevant memories from the step $\tau{-}1$ vector database, concatenated with the new session $s_{\tau}$, mirroring the inference-time setup.
To enable dense session-level rewards, we generate 5 QA pairs per session (12,000 pairs total) using Claude 4.5 Opus. The generation model receives the new session along with the retrieved memories, allowing it to create questions that test both current session retention and connections to prior history. Questions span three types: single-session (testing current session content), multi-session (requiring cross-session aggregation), and temporal-reasoning (involving time-based inference).
Reward Computation.
During GRPO training (8 rollouts per session, 5 epochs), each rollout $i$ at session $\tau$ produces candidate memory bank $M^{(i)}_{\tau}$. GPT-4.1-mini answers the pre-generated questions by retrieving from $M^{(i)}_{\tau}$, and an LLM judge (GPT-4.1-mini) evaluates correctness against ground-truth answers. The average accuracy across 5 questions yields the task reward $r_{\textrm{task}}$ (Eq. 6). Within each rollout, all four memory types share this reward, differentiated through contribution-aware gradient weighting (Section 3.4.2). This produces 96,000 session-rollouts for policy optimization. A human evaluation on 100 sampled QA pairs confirms that 92% of generated answers are factually correct (Appendix H.2).
@@ 4.4.1 Effect of Gradient Weighting
To investigate the effect of contribution-aware gradient weighting (Section 3.4.2), we vary the weighting coefficient $\alpha$ that amplifies updates for the dominant contributing memory type. We conduct this ablation on a reduced training set for efficiency. As shown in Figure 4, performance improves as $\alpha$ increases from 1 (no weighting, 82.82%) to 4 (84.04%), confirming that attributing credit to high-contribution memory types enhances the final model performance. However, excessively large $\alpha$ values degrade performance due to gradient imbalance among memory types, with the optimal value at $\alpha=4$. A human evaluation further confirms that retrieval frequency is a reliable attribution proxy, with the dominant memory type aligning with the actual information source in 95% of cases (Appendix H.1).
@@ 4.4.3 Answer Model Generalization
To evaluate whether the constructed memory generalizes across different answer models, we fix the memory construction model to our trained Qwen3-4B and vary the answer model. Table 2 shows that our memory maintains high quality across answer models of varying capabilities. We further investigate the feasibility of replacing closed-source answer models with open-source alternatives in Appendix H.3, where results show that Qwen3-30B-A3B achieves performance comparable to GPT-4.1-mini on LoCoMo when evaluated on the same high-quality memory.
@@ A.1 Benchmark Datasets
Appendix B Training Details
B.1 SFT Training